[figure: 9.7 cm]

EFFECT OF PRESSURE ON MAGNETIC TRANSITIONS IN PrMnGe

R. Duraj^{a} and A. Szytuła^{b} aInstitute of Physics, Technical University of Cracow bInstitute of Physics, Jagiellonian University, Cracow

The compund PrMnGe crystallizes in the tetragonal CeFeSi-type structure. Pr, Mn and Ge atoms are arranged in alternate layers stacked along the c axis in the sequence PrGeMn_{2}GePrPrGeMn_{2}GePr. The compound is antiferromagnetic below T_{N}=410K. At T_{C}=150K it exhibits ferromagnetic behaviour followed by a ferro-antiferromagnetic transition at temperature T_{M} below 80K, simultaneously with a crystallographic phase transition (tetragonal to orthorhombic symmetry). The influence of pressure on T_{C} and T_{M} have been investigated by means of ac susceptibility measurements under hydrostatic pressure. The magnetic (p,T) phase diagram is presented. T_{C} decrease slightly with increasing pressure. On the other hand, a positive effect of pressure is observed for T_{M}. At pressure above 1.5GPa (∼ 1.7GPa) a multicritical point is observed. The results are discussed in comparison to those obtained for compounds RT_{2}X_{2} in the ThCr_{2}Si_{2}-type structure.

[figure: 13.4 cm]

Subject category :

5. Phase Transitions and Critical Phenomena

Presentation mode : poster

Corresponding author : R.Duraj

Address for correspondence : Institute of Physics, Technical University of Cracow ul. Podchora̧żych 1 30-084 Cracow

Email address : puduraj@cyf-kr.edu.pl